PI: Guirado
Title: EQPegB

Session: Oct 1996 (session 3)
Frequency: 3.6 cm
Mode: A
Obs time (UT): 302 EU 15:30(28/10)-01:30(29/10)
Stations: EbNtOn
Processor: Bonn
EU Mk3 Tapes/St.: 4


To give feedback on this experiment, click here

Experiment Feedback

Onsala (J. Conway, A. Polatidis at Wed Oct 30 03:57:23 EST 1996) Success
No known problems.

Noto (Corrado Trigilio at Tue Nov 5 11:24:40 EST 1996) Success
No problems reported. Good sky conditions.

Effelsberg (D Graham at Wed Nov 6 09:06:54 EST 1996) Minor failures
1554 UT to 1643 no source change, communication problem

MPIfR (Scott Aaron at Tue Sep 23 14:30:50 METDST 1997) Success
No problems.